OpenStudy (anonymous):

**WILL MEDAL** Janine inflated 1 ball to a radius of 18 cm and another ball to a radius of 12 cm. How much more air was in the larger ball? Use 3.14 to approximate pi and express your answer in hundredths. ___cm3

OpenStudy (anonymous):

well the volume of a sphere is 4/3pir^3 so you would take the volume of both and then take the difference

OpenStudy (anonymous):

So 129.5567

OpenStudy (isaiah.feynman):

Difference in volume is given by \[V-v = \frac{ 4 }{ 3 }\pi R^{3} - \frac{ 4 }{3 } \pi r^{3}\]

OpenStudy (anonymous):

well i got a different answer i did 4/3pi(18)^3 - 4/3pi(12)^3 is this what you did

OpenStudy (anonymous):

Im confused

OpenStudy (isaiah.feynman):

We can then simplify by doing, \[V-v = \frac{ 4 }{ 3 }\pi(R^{3}-r^{3})\]

OpenStudy (isaiah.feynman):

Where big r is the radius of the big ball and small r is the radius of the small ball.

OpenStudy (anonymous):

Im still confused whats the answer?

OpenStudy (anonymous):

well just take it step by step 18^3-12^3 is what

OpenStudy (anonymous):

5832-1728

OpenStudy (anonymous):

and that is 4104

OpenStudy (anonymous):

the just multiply that by 4 then divide by 3

OpenStudy (anonymous):

after that multiply by pi

OpenStudy (anonymous):

17182.08?

OpenStudy (anonymous):

@MelodicSymphony

OpenStudy (anonymous):

haha i think so

OpenStudy (anonymous):

just checked you are right

OpenStudy (anonymous):

Thx ill tell u in like 3 mins what i get